OpenStudy (anonymous):

Find the six trigonometric functions of theta= 450 degrees

OpenStudy (anonymous):

check your unit circle at 90 degrees

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

I really don't know.How to answer this.I have no idea on what to do and how to solve this.Please help me. I will pass this for my next subject.

OpenStudy (anonymous):

sin (450º) = sin (90º) = sin (270º) cos (450º) = cos (90º) = cos (270º) tan (450º) = tan (90º) = tan (270º) cot (450º) = cot (90º) = cot (270º) sec (450º) = sec (90º) = sec (270º) csc (450º) = csc (90º) = csc (270º)

OpenStudy (anonymous):

Is this the answer? It must be in reciprocal.

OpenStudy (anonymous):

sin theta =opposite over hypotenuse

OpenStudy (anonymous):

sin 450= 1 cos 450= 0 tan 450= undefined csc 450= 1 sec 450= undefined cot 450= 0

OpenStudy (anonymous):

and cosine theta = adjacent over hypotenuse tan theta =opposite over adjacent

OpenStudy (anonymous):

I know but is this for a triangle or just for the 450 degrees?

OpenStudy (anonymous):

for a triangle.

OpenStudy (anonymous):

can you post a pic of it?

OpenStudy (anonymous):

Please answer this for me.Thanks and God bless you.

OpenStudy (anonymous):

Find the six trigonometric functions of theta = -270 degrees Find the six trigonometric functions of theta = 45 degrees remember this is for a triangle.

OpenStudy (anonymous):

Time to go. I have my class now. TRIGONOMETRY.Thanks anyway. GOD BLESS YOU! and take care.Continue to hep others.

OpenStudy (anonymous):

will do and i answered your first question

OpenStudy (anonymous):

Thank you very much for your help today.Your answer in question number 1 and two are both correct. You really good at this huh. Keep up the good work.GOD BLESS YOU!

OpenStudy (anonymous):

DiotropicInventor, can I ask a favor again? Please help me.Find the six trigonometric funtions of 1.)30 degrees, 2.) -45 degrees 3.) 135 degrees 4.) 240 degrees, 5. -33o degrees. And can you please draw it in a standard position? please. and also can you tell me how do you do it? please. GOD BLESS. I need to know it because next week is our MIDterm exam.

OpenStudy (anonymous):

in number 5. what I mean is -330...thanks..